OpenStudy (anonymous):

Solve (x-6)-(2x-10)=2

OpenStudy (anonymous):

expand brackets. x-6-2x+20 = 2 collect like terms -x + 14 = 2 -x = -12 x = ??? (you can do this)

OpenStudy (anonymous):

i got .5

OpenStudy (anonymous):

@marixoxo76 do u get it?

OpenStudy (anonymous):

and it's not 0.5

OpenStudy (anonymous):

yes i got it it's 6 thanks

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

@marixoxo76 , 6 is wrong.

OpenStudy (anonymous):

for these type of questions you can check if you are right by substituting the value u found in the end. You said x=6 but u get -2 not 2 so it is wrong.

OpenStudy (anonymous):

jay when ya 1st put the equation in you havex-6-2x+20 = 2 but how did you get +20?

OpenStudy (anonymous):

i got +10

OpenStudy (anonymous):

oh woops sorry, I thought I saw a 2 there, I'm going blind.

OpenStudy (anonymous):

expand brackets. x-6-2x+10 = 2 collect like terms -x + 4 = 2 -x = -2 x = ??? (you can do this)

OpenStudy (anonymous):

its .5 isnt it?

OpenStudy (anonymous):

yea im pretty sure its .5

OpenStudy (anonymous):

I told u already that is wrong and you can test it urself by substituting that value of x=0.5 into the original equation.

OpenStudy (anonymous):

oh my bad i was doing vice versa its 2